Solve the system of equation by the substitution method
x+y=20
5x+y=0

Answer by algebrahouse.com(1659)

x + y = 20
5x + y = 0 ----> y = -5x {subtracted 5x from each side}

x - 5x = 20 {substituted -5x, in for y, into first equation}
-4x = 20 {combined like terms}
x = -5 {divided each side by -4}

x + y = 20 {first equation}
-5 + y = 20 {substituted -5, in for x, into first equation}
y = 25 {added 5 to each side}

x = -5 and y = 25

For more help from me, visit: www.algebrahouse.com